OpenStudy (anonymous):

Find the inverse of each function. y=3log4x^2 Step by step please. I'm stupid ..

OpenStudy (anonymous):

i'm stupid too...:(

OpenStudy (anonymous):

Is it \[ y= 3 \ln ( 4x^2) \]

OpenStudy (anonymous):

\[ y =6 \ln (2 x)\\ \frac y 6 = \ln (2 x)\\ e^{\frac y 6}= e^{\ln 2x}=2x\\ x= \frac 12 e^{\frac y 6} \] the inverse is \[ g(x)= \frac 12 e^{\frac x 6} \]